OpenStudy (anonymous):

please help, solve the equation, round to the nearest tenth if necessary. 7r-7=2r+18

OpenStudy (anonymous):

7r-2r=18+7 5r=25 r=25/5 r=5

OpenStudy (anonymous):

That is the idea I was getting but I needed extra help Thank You so much for your help !!!

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

one more do you think you could help @adamarbour

OpenStudy (anonymous):

2x+12=18-x

OpenStudy (ccswims):

add like terms, like above

OpenStudy (anonymous):

i tried then i got 2x-x and i got lost\

OpenStudy (ccswims):